# BayesianFitForecast: A User-Friendly R Toolbox for Parameter Estimation and Forecasting with Ordinary Differential Equations

### v2

@@ Abstract
Mathematical models based on ordinary differential equations (ODEs) are essential tools across various scientific disciplines, including biology, ecology, and healthcare informatics where they are used to simulate complex dynamic systems and inform decision-making. In this paper, we introduce BayesianFitForecast, an R toolbox specifically developed to streamline Bayesian parameter estimation and forecasting in ODE models, making it particularly relevant to health informatics and public health decision-making (https://github.com/gchowell/BayesianFitForecast/). This toolbox enables automatic generation of Stan files, allowing users to configure models, define priors, and analyze results with minimal programming expertise. To demonstrate the versatility and robustness of BayesianFitForecast, we apply it to the analysis of the 1918 influenza pandemic in San Francisco, comparing Poisson and negative binomial error structures within the SEIR model. BayesianFitForecast provides robust tools for evaluating model performance, including convergence diagnostics, posterior distributions, credible intervals, and performance metrics. By improving the accessibility of advanced Bayesian methods, this toolbox significantly broadens the application of Bayesian inference methods to dynamical systems critical for healthcare and epidemiological forecasting. Additionally, we include a tutorial video demonstrating the toolbox’s functionality (https://youtu.be/jnxMjz3V3n8).
Keywords: Bayesian workflow, real-time forecasting, performance, epidemiology, healthcare informatics
@@ 1   Background
However, implementing Bayesian inference in ODE models often requires extensive programming knowledge and manual coding, presenting a significant barrier for many researchers. Tools like Stan [16] facilitate the implementation of Bayesian estimation and forecasting, allowing for rigorous uncertainty quantification and model validation [17, 18, 19, 20, 21, 22, 23]. Yet, the technical complexity involved in writing code and the need for specialized programming skills can deter many potential users.
In response to these challenges, we introduce BayesianFitForecast, a user-friendly R toolbox specifically designed to streamline the Bayesian estimation process, making it accessible to a broad audience, including graduate students in applied mathematical and statistical sciences. Traditionally, employing Bayesian estimation necessitates writing code in Stan [17, 4], a probabilistic programming language. However, our toolbox eliminates this barrier by automatically generating the necessary Stan files based on user-defined options. BayesianFitForecast offers robust tools for evaluating model performance, including convergence diagnostics, posterior distributions, credible intervals, and performance metrics, all with minimal programming expertise required. We have included a detailed tutorial within this paper to guide users through the various features and functionalities of the toolbox, ensuring that even those with minimal programming experience can effectively leverage Bayesian methods with ODE models. A tutorial video demonstrating the toolbox’s functionality is available on YouTube (https://youtu.be/jnxMjz3V3n8).
@@ 2.2.1 Bayesian inference
  2. 2.
    b) Uniform priors. When using uniform priors, we assume minimal prior knowledge about the parameters. This involves establishing a range for parameter values without assuming they center around any specific value. Typically, a uniform distribution is used for this purpose. Uniform priors are particularly useful when reliable prior information about the parameters is lacking. By using uniform distributions, we allow the data to play a more significant role in shaping the posterior distributions of the parameters. This method is especially advantageous in the early stages of an outbreak when specific prior knowledge may not be available.
The error structures, such as the negative binomial distribution and normal distribution, have additional variance-related parameters, whose prior distributions will be specified in the toolbox illustration.
@@ 2.2.2 Numerical Implementation
Stan and MCMC. Stan is a probabilistic programming language designed for Bayesian statistical modeling and inference, utilizing the MCMC algorithm to sample from complex posterior distributions [4]. Stan employs a sophisticated variant of Markov chain Monte Carlo (MCMC) called Hamiltonian Monte Carlo (HMC) [24], which leverages gradient information for efficient sampling, and the No-U-Turn Sampler (NUTS), an adaptive algorithm that automatically adjusts the trajectory length to avoid revisiting previous states. This approach enhances convergence speed and sampling accuracy without manual tuning. Stan’s automatic differentiation capability ensures precise computation of gradients, which is crucial for HMC performance. By using Stan, users can specify a wide variety of statistical models and obtain samples from the posterior distribution, facilitating parameter estimation, uncertainty quantification, and predictive inference. In this toolbox, the Stan file is generated automatically for parameter estimation, eliminating the need for users to write Stan code manually.
@@
Model Selection and Quality of Model Fit. To compare the quality of model fit, we use the Deviance Information Criterion (DIC), a Bayesian generalization of the Akaike Information Criterion (AIC). The DIC is given by:
$\text{DIC}=\bar{D}+p_{D}$ (5)
where $\bar{D}$ is the mean deviance and $p_{D}$ is the effective number of parameters. This metric accounts for model complexity regarding the number of model parameters and is used for model selection in the Bayesian framework.
@@ 2.3 Performance metrics
The weighted interval score (WIS) [25, 27], is a proper score that provides quantiles of predictive forecast distribution by combining a set of Interval Scores (IS) for probabilistic forecasts. An IS is a simple proper score that requires only a central $(1-\alpha)\times 100\%PI$ [25] and is described as:
@@
  • The sharpness of $F$, given by the width $u-l$ of the central $(1-\alpha)\times 100\%PI$.
@@ 2.4.1 Installation and setup
  • Ensure that Rtools and Java are downloaded and up to date. You can download Rtools from here and Java from here.
@@
  • If running the code in RStudio for the first time, it is necessary to install all required packages before proceeding. To do so, navigate to the Packages tab in the upper-right panel of RStudio, and select Install. In the dialog box, enter the names of the relevant packages: bayesplot, readxl, xlsx, openxlsx, rstan, and ggplot2.
@@ 2.4.2 File descriptions
Upon downloading the relevant folder, you will find several files inside. Some of these files are meant to be used, while others should remain untouched. The untouched files are functions utilized in other files. Table 1 provides an overview of each file in the folder and its respective function or purpose. The files that the user should work with are: "options_generaluse.R", "run_MCMC.R" and "run_analyzeResults.R".
@@
where $\beta>0$ represents the transmission rate, $\kappa>0$ the incubation rate, $\gamma>0$ the recovery rate, $\rho\in[0,1]$ the reporting proportion, and $N$ the total population size which is assumed to be known. A compartmental diagram of this model is given in Figure 2. In this model, $S$, $E$, $I$, $R$, and $C$ are called the state variables and $\beta$, $\kappa$, $\gamma$, $\rho$, and $N$ are called parameters. When defining the variables ("vars"), they should be listed in the same order as their derivatives appear in the system of ODEs. Please note that each variable should be defined as a string (enclosed in double quotes). Therefore, for this model, we have:
@@
$S$$E$$I$$R$Observations are new cases; $y(t)=\kappa\rho E(t)$$\frac{\beta SI}{N}$$\kappa E$$\gamma I$
@@
As another example, we can introduce the SEIRD model (Figure 3). In addition to the states $S$, $E$, $I$, and $R$, the SEIRD model also tracks the number of disease-induced deaths (denoted as $D$) and is given by:
@@
$S$$E$$I$$R$$D$Observations are new deaths; $y(t)=\gamma\rho I(t)$$\frac{\beta SI}{N}$$\kappa E$$\gamma(1-\rho)I$$\gamma\rho I$
@@
$S$$E$$I$$U$$R$Observations are new cases; $y(t)=\kappa\rho E(t)$$\frac{\beta(I+U)S}{N}$$\kappa\rho E$$\kappa(1-\rho)E$$\gamma I$$\gamma U$
@@ 2.4.19 Prior distribution for additional parameters in error structure
    "Ic": Specify the initial condition as a string. For example, in the SEIR model, if the new parameter is named $i_{0}$, then set Ic = ("N-$i_{0}$", 0,"$i_{0}$", 0, "$i_{0}$"). Ensure all letters are enclosed in double quotation marks (strings).
@@ 2.4.26 Saved data
  • left=2em
    “model_name”
  • left=2em
    “calibrationperiod”
  • left=2em
    “forecastinghorizon”
  • left=2em
    “errstrc”
  • left=2em
    “caddisease”
@@
By running this file, the user can generate and analyze the results. It is important to keep the option file the same as when running "run_MCMC"(except the forecasting horizon to achieve a zoomed version as already explained). Additionally, line 10 of this file needs to be modified to include the name of the option file you are using. After running this code, several results will be generated in the output folder. The files are formatted as follows:
@@
  • Convergence: This file includes the convergence parameter "Rhat". It is recommended to use the results obtained by Bayesian inference when the parameter "Rhat" is less than 1.1. To achieve this goal, you can increase "niter", the number of iterations, or "num_chain", the number of chains. To illustrate the format of this file, consider the example: convergence-Bayesian-weak-
    sanfrancisco-normal-cal-10-fcst-10
  • Forecast: This file includes five columns: Date, Data, Median, Lower Bound and Upper Bound of the 95% PIs. It can be used to plot the fit easily using the given values. To illustrate the format of this file, consider the example: "forecast-Bayesian-weak-
    sanfrancisco-normal-cal-10-fcst-10"
@@
  • Performance Metrics: This file provides four performance metrics for both calibration and forecasting periods: MAE, MSE, WIS, and Coverage of the 95% PIs. It is important to note that if data is unavailable for the forecasting period, the performance metrics will show NA, which is not applicable for the forecasting part. To illustrate the format of this file, consider the example: "performance metrics-Bayesian-weak-sanfrancisco-
    normal-cal-10-fcst-10"
@@
  • Forecast: The plot for fitting the model and forecasting will be found here. It shows the median fit, the margins showing the 95%PI, the available data points, and the calibration line. To illustrate the format of this file, consider the example: "Forecast-Bayesian-weak
    -sanfrancisco-normal-cal-10-fcst-10"
  • Trace plot: A trace plot is a graphical representation of the sampled values of a parameter over iterations of the MCMC algorithm. It shows how the parameter values change from one iteration to the next, allowing you to visually inspect the behavior of the sampler. To illustrate the format of this file, consider the example: "traceplot-Bayesian-weak-
    sanfrancisco-normal-cal-10-fcst-10"

### v3

@@ Abstract
Background: Mathematical models based on ordinary differential equations (ODEs) are essential tools across various scientific disciplines, including biology, ecology, epidemic modeling, and healthcare informatics, where they are used to simulate complex dynamic systems and inform decision-making. However, implementing Bayesian calibration and forecasting typically requires substantial coding in Stan or similar tools. To support Bayesian parameter estimation and forecasting for such systems, we introduce BayesianFitForecast, a user-friendly R toolbox specifically developed to streamline Bayesian parameter estimation and forecasting in ODE models, making it particularly relevant to health informatics and public health decision-making (https://github.com/gchowell/BayesianFitForecast/).
Results: This toolbox enables automatic generation of Stan files, allowing users to configure models, define priors, and analyze results with minimal programming expertise. By eliminating manual coding, BayesianFitForecast significantly lowers the technical barrier to Bayesian inference with dynamical systems. We demonstrate its flexibility and usability through applications to historical epidemic datasets (e.g., the 1918 influenza pandemic in San Francisco and the 1896–1897 Bombay plague) and simulated data, showing robust parameter estimation and forecasting performance under Poisson and negative binomial observation error structures. The toolbox also provides robust tools for evaluating model performance, including convergence diagnostics, posterior distributions, credible intervals, and performance metrics.
Conclusion: By improving the accessibility of advanced Bayesian methods, BayesianFitForecast broadens the application of Bayesian inference in time-series modeling, healthcare forecasting, and epidemiological applications. In addition to the R scripting interface, a built-in Shiny web application is included, enabling interactive model configuration, visualization, and forecasting. A tutorial video demonstrating the toolbox’s functionality is also available (https://youtu.be/jnxMjz3V3n8).
keywords:
Bayesian inference, Bayesian calibration, MCMC sampling, Hamiltonian Monte Carlo, Forecasting, Model selection, Uncertainty quantification, Epidemiological modeling, ODE parameter estimation, Stan R interface.
@@ 1 Background
However, implementing Bayesian calibration and forecasting for ODE-based systems often requires extensive programming knowledge and manual coding, presenting a significant barrier for many researchers. Tools like Stan [16] facilitate the implementation of Bayesian estimation and forecasting, allowing for rigorous uncertainty quantification and model validation [17, 18, 19, 20, 21, 22, 23]. Yet, the technical complexity involved in writing code and the need for specialized programming skills can deter many potential users.
In response to these challenges, we introduce BayesianFitForecast, a user-friendly R toolbox specifically designed to streamline the Bayesian calibration and forecasting process for real-world epidemic data and other dynamic systems, making it accessible to a broad audience, including graduate students in applied mathematical and statistical sciences. Traditionally, employing Bayesian estimation necessitates writing code in Stan [17, 4], a probabilistic programming language. However, our toolbox eliminates this barrier by automatically generating the necessary Stan files based on user-defined options. BayesianFitForecast offers robust tools for evaluating model performance, including convergence diagnostics, posterior distributions, credible intervals, and performance metrics, all with minimal programming expertise required. We have included a detailed tutorial within this paper to guide users through the various features and functionalities of the toolbox, ensuring that even those with minimal programming experience can effectively leverage Bayesian methods with ODE models. The toolbox and accompanying tutorial materials are available at (https://github.com/gchowell/BayesianFitForecast) with a full video walkthrough at (https://youtu.be/jnxMjz3V3n8).
@@ 2.2.1 Bayesian inference
b) Uniform priors. When using uniform priors, we assume minimal prior knowledge about the parameters. This involves establishing a range for parameter values without assuming they center around any specific value. Typically, a uniform distribution is used for this purpose. Uniform priors are particularly useful when reliable prior information about the parameters is lacking. By using uniform distributions, we allow the data to play a more significant role in shaping the posterior distributions of the parameters. This method is especially advantageous in the early stages of an outbreak when specific prior knowledge may not be available. The error structures, such as the negative binomial distribution and normal distribution, have additional variance-related parameters, whose prior distributions will be specified in the toolbox illustration.
@@ 2.2.2 Numerical Implementation
Stan and MCMC. Stan is a probabilistic programming language designed for Bayesian statistical modeling and inference, utilizing the MCMC algorithm and the No-U-Turn Sampler (NUTS) to sample from complex posterior distributions [4]. Stan employs a sophisticated variant of Markov chain Monte Carlo (MCMC) called Hamiltonian Monte Carlo (HMC) [24], which leverages gradient information for efficient sampling, and the No-U-Turn Sampler (NUTS), an adaptive algorithm that automatically adjusts the trajectory length to avoid revisiting previous states. This approach enhances convergence speed and sampling accuracy without manual tuning. Stan’s automatic differentiation capability ensures precise computation of gradients, which is crucial for HMC performance. By using Stan, users can specify a wide variety of statistical models and obtain samples from the posterior distribution, facilitating Bayesian calibration, uncertainty quantification, and predictive inference, especially in forecasting ODE-based systems driven by real-world epidemic data. In this toolbox, the Stan file is generated automatically for parameter estimation, eliminating the need for users to write Stan code manually.
@@
Model Selection and Quality of Model Fit. To compare the quality of model fit, we use several criteria that are widely adopted in Bayesian modeling, including the Deviance Information Criterion (DIC), Watanabe-Akaike Information Criterion (WAIC), and Leave-One-Out cross-validation (LOO) [25, 26, 27].
The DIC is a Bayesian generalization of the Akaike Information Criterion (AIC) and is given by:
$\text{DIC}=\bar{D}+p_{D}$
where $\bar{D}$ is the mean deviance and $p_{D}$ is the effective number of parameters. DIC provides a balance between model fit and complexity, penalizing over-parameterization.
The WAIC is a fully Bayesian alternative that approximates Bayesian cross-validation. It is calculated using the log pointwise predictive density (lppd) and includes a penalty term for model complexity based on the variance of the log-likelihood:
$\text{WAIC}=-2\left(\text{lppd}-p_{\text{WAIC}}\right)$
where $p_{\text{WAIC}}$ is the estimated effective number of parameters. WAIC is asymptotically equivalent to Bayesian cross-validation and is especially useful for hierarchical models. BayesianFitForecast calculates WAIC automatically from the posterior samples, enabling users to compare models with different structures or prior assumptions based on out-of-sample predictive accuracy.
Lastly, LOO uses leave-one-out cross-validation by approximating the model’s predictive accuracy when each data point is held out in turn. We employ Pareto-smoothed importance sampling (PSIS-LOO), a computationally efficient approach to estimate LOO from a single posterior sample. LOO provides a robust, data-driven estimate of out-of-sample predictive performance, making it a valuable tool for model comparison, particularly when overfitting is a concern.
@@ 2.3 Performance metrics
The weighted interval score (WIS) [28], is a proper score that provides quantiles of predictive forecast distribution by combining a set of Interval Scores (IS) for probabilistic forecasts. An IS is a simple proper score that requires only a central $(1-\alpha)\times 100\%PI$ [28] and is described as:
@@
  • The sharpness of $F$, given by the width ulu-l of the central $(1-\alpha)\times 100\%PI$.
@@ 2.4.1 Installation and Setup
  • Install R (recommended version <= 4.5) from the Comprehensive R Archive Network (CRAN) and RStudio as the integrated development environment.
  • Ensure that Rtools and Java are downloaded and up to date.
@@
  • If running the code in RStudio for the first time, it is necessary to install all required packages before proceeding. To do so, navigate to the Packages tab in the upper-right panel of RStudio, and select Install. In the dialog box, enter the names of the following packages: bayesplot, readxl, xlsx, openxlsx, rstan, ggplot2, loo, stringr, and gridExtra.
  • Important: This toolbox is compatible with R versions prior to 4.5 and Stan versions prior to 2.36. Using later versions may lead to installation or runtime errors due to major changes in package dependencies and Stan behavior.
@@ 2.4.2 File descriptions
Upon downloading the relevant folder, you will find several files inside. Some of these files are meant to be used, while others should remain untouched. The untouched files are functions utilized in other files. Table 1 provides an overview of each file in the folder and its respective function or purpose. The main files that the user should work with are: "options_generaluse.R", "run_MCMC.R" and "run_analyzeResults.R". In addition, the user can generate the prior solution of the model by running prior_sltn.R.
@@
prior_sltn.R This file generates the forward solution of the model by sampling from the prior distribution.
@@
where $\beta>0$ represents the transmission rate, $\kappa>0$ the incubation rate, $\gamma>0$ the recovery rate, $\rho\in[0,1]$ the reporting proportion, and $N$ the total population size which is assumed to be known. A compartmental diagram of this model is given in Figure 2. In this model, $S$, $E$, II, $R$, and $C$ are called the state variables and $\beta$, $\kappa$, $\gamma$, $\rho$, and $N$ are called parameters. When defining the variables ("vars"), they should be listed in the same order as their derivatives appear in the system of ODEs. Please note that each variable should be defined as a string (enclosed in double quotes). Therefore, for this model, we have:
@@
SSEEIIRRObservations are new cases; $y(t)=\kappa\rho E(t)$$\frac{\beta SI}{N}$$\kappa E$$\gamma I$
@@
As another example, we can introduce the SEIRD model (Figure 3). In addition to the states $S$, $E$, II, and $R$, the SEIRD model also tracks the number of disease-induced deaths (denoted as $D$) and is given by:
@@
SSEEIIRRDDObservations are new deaths; $y(t)=\gamma\rho I(t)$$\frac{\beta SI}{N}$$\kappa E$$\gamma(1-\rho)I$$\gamma\rho I$
@@
SSEEIIUURRObservations are new cases; $y(t)=\kappa\rho E(t)$$\frac{\beta(I+U)S}{N}$$\kappa\rho E$$\kappa(1-\rho)E$$\gamma I$$\gamma U$
@@
2.4.9 time_dependent_templates
The time_dependent_templates object defines a set of time-dependent parameters, which are useful when the ODE system includes parameters that vary with time.
To incorporate such parameters into the model, simply define them as string functions, as shown in the examples below. Each function should be expressed in terms of the time variable t and elements from the params vector. These strings can later be parsed and injected into the model code (see 3.7).
Here are some common examples:
            time_dependent_param1 <- "            return (params1 * exp(-params2 * t));            "
            time_dependent_param2 <- "            if (t < params1) {              return(params2);            } else if (t < params3) {              return(params2 + (params4 - params2) * (t - params1)              / (params3 - params1));            } else {              return(params4);            }            "
            time_dependent_param3 <- "            if (t < params1) {              return(params2);            } else {              return(params3);            }            "
            time_dependent_param4 <- "            return (params1 + params2 * sin(2 * pi() * t / params3));            "
            time_dependent_param5 <- "            return (params1 * exp(-0.5 * square((t - params2) / params3)));            "In summary, a time-dependent parameter is simply a user-defined function of time and model parameters. These functions reference the parameter values already included in the params vector and are designed to be evaluated at runtime within the ODE solver or model function.
@@
If the ODE system includes any time-dependent parameter, simply write it as time_dependent_param1 as defined in the time_dependent_templates list.
params <- c("beta0","beta1", "q", "kappa", "gamma", "rho", "N", "t_int")time_dependent_templates <- list(  time_dependent_param1 = "if (t < params8) { return (params1); }  else { return (params2 + (params1 - params2)  * exp(-params3 * (t - params8))); }")ode_system <- ’  diff_var1 = -time_dependent_param1 * vars1 * vars3 / params7  diff_var2 = time_dependent_param1 * vars1 * vars3 / params7                                    - params4 * vars2  diff_var3 = params4 * vars2 - params5 * vars3  diff_var4 = params5 * vars3  diff_var5 = params4 * params6 * vars2’
@@ 2.4.20 Prior distribution for additional parameters in error structure
    "Ic": Specify the initial condition as a string. For example, in the SEIR model, if the new parameter is named $i_{0}$, then set Ic = ("N-$i_{0}$", $i_{0}", 0$, $i_{0}$"). Ensure all letters are enclosed in double quotation marks (strings).
@@
2.4.24 prior_sltn.R
prior_sltn.R performs prior predictive simulations by sampling parameter values from user-defined prior distributions and using them to solve the specified ODE model. For each draw from the prior, the function integrates the system of equations to produce a trajectory, and the collection of these trajectories forms an ensemble that reflects the uncertainty inherent in the prior assumptions. To run the function, the user must define the ODE model, including its parameters and state variables, and specify the prior distributions for the parameters in the options file. It is worth noting that the main functions of this toolbox, run_MCMC and run_analyzeResults, which generate the posterior distribution of the parameters, will be described later. The prior_sltn.R function is used solely to generate the prior solution of the model using the prior distribution of parameters.
As an example, we demonstrate the use of the prior_sltn.R function with the SEIR model described earlier. Simulated data for the incidence cases ($\kappa E$) are generated by solving the SEIR model forward using the parameter values $\beta=0.6$, $\gamma=0.1$, $\kappa=0.2$, $\rho=0.9$, and N=550, [^0] N=550{,}000, and are stored in SEIRC.xlsx. To reflect strong prior knowledge, we specify informative normal prior distributions centered at these true values, each with a small standard deviation ($\sigma=0.01$): beta <- "normal(0.6, 0.01)T[0,]", gamma <- "normal(0.1, 0.01)T[0,]", kappa <- "normal(0.2, 0.01)T[0,]", and rho <- "normal(0.9, 0.01)T[0,]". These specifications, defined in options_SEIR_sanfrancisco_prior.R, constrain the sampled parameter values to remain very close to the truth, resulting in simulated trajectories that closely align with the simulated dataset.
In order to use the prior_sltn.R function, we need to update line 10 of the code to include the option file options_SEIR_sanfrancisco_prior.R.
Figure 5 shows prior uncertainty (95% CI) for $\dfrac{dC}{dt}$ and the resulting ODE state variable trajectories. The tight prior distribution for the parameters yields low variability in the simulated trajectories, representing our strong initial beliefs before observing data.
Figure 5: Prior predictive simulation results using strong informative priors: (a) incidence cases and (b) all state variables.
@@ 2.4.28 Saved data
  • [left=2em]
  • “model_name"
  • “calibrationperiod"
  • “forecastinghorizon"
  • “errstrc"
  • “caddisease"
@@
By running this file, the user can generate and analyze the results. It is important to keep the option file the same as when running "run_MCMC"(except the forecasting horizon to achieve a zoomed version as already explained). Additionally, line 13 of this file needs to be modified to include the name of the option file you are using. After running this code, several results will be generated in the output folder. The files are formatted as follows:
@@
  • [left=2em]
  • Convergence: This file includes the convergence parameter "Rhat". It is recommended to use the results obtained by Bayesian inference when the parameter "Rhat" is less than 1.1. To achieve this goal, you can increase "niter", the number of iterations, or "num_chain", the number of chains. To illustrate the format of this file, consider the example: convergence-Bayesian-weak-sanfrancisco-normal-cal-10-fcst-10.
  • Forecast: This file includes five columns: Date, Data, Median, Lower Bound and Upper Bound of the 95% PIs. It can be used to plot the fit easily using the given values. To illustrate the format of this file, consider the example: "forecast-Bayesian-weak-sanfrancisco-normal-cal-10-fcst-10".
@@
  • Performance Metrics: This file provides four performance metrics for both calibration and forecasting periods: MAE, MSE, WIS, and Coverage of the 95% PIs. It is important to note that if data is unavailable for the forecasting period, the performance metrics will show NA, which is not applicable for the forecasting part. To illustrate the format of this file, consider the example: "performance metrics-Bayesian-weak-sanfrancisco-normal-cal-10-fcst-10".
@@
  • Forecast: The plot for fitting the model and forecasting will be found here. It shows the median fit, the margins showing the 95%PI, the available data points, and the calibration line. To illustrate the format of this file, consider the example: "Forecast-Bayesian-weak-sanfrancisco-normal-
    cal-10-fcst-10"
  • Trace plot: A trace plot is a graphical representation of the sampled values of a parameter over iterations of the MCMC algorithm. It shows how the parameter values change from one iteration to the next, allowing you to visually inspect the behavior of the sampler. To illustrate the format of this file, consider the example: "traceplot-Bayesian-weak-sanfrancisco-normal-cal-10-fcst-10".